A stamping plant supplies sheet metal parts to a final assembly plant in the automotive industry. Annual demand for a typical part is 150,000 pc. Average cost per piece is $20; holding cost is 25%; changeover (setup) time for the presses is 5 hr; and cost of downtime for changing over a press is $200/hr. Compute the economic batch size and the total annual inventory cost for the data.
Answer #1

Let us first calculate the cost of setup (C).

For that, we have Changeover (setup) time for the presses is = 5 hr

Cost of downtime is $200/hr

C = 200*5 = $1000

Now, Let's calculate Economic Batch Size.

For that we have Annual demand (A) = 150,000 pc

Average Cost per piece = $20

Holding Cost = 25%

So, holding cost = 20*25%

= (20)0.25

= $5

Economic Batch Size =

So, the Economic Batch Size is 7746 pieces.

Now let's calculate Total annual inventory cost = (7746/2)*(5) + (150000/7746)*(1000)

= 19365 + 19365

= $38730